OpenStudy (anonymous):

Five hundred tickets were sold for a certain music concert. The tickets for the adults and children sold for $7.50 and $4.00, respectively, and the total receipts for the performance were $3,312.50. How many of each kind of ticket were sold? The expression could be 7.50x+4.00y=3,312.50

OpenStudy (anonymous):

find the answer

OpenStudy (anonymous):

show me how to solve please

OpenStudy (anonymous):



OpenStudy (texaschic101):

x + y = 500 ---> x = 500 - y 7.50x + 4y = 3312.50 7.50(500 - y) + 4y = 3312.50 3750 - 7.50y + 4y = 3312.50 -- subtract 3750 from both sides -7.50y + 4y = 3312.50 - 3750 -- combine like terms -3.50y = - 437.50 -- divide both sides by -3.50 y = 437.50/3.50 y = 125 x + y = 500 x + 125 = 500 x = 500 - 125 x = 375 check.. 7.5(375) + 4(125) = 3312.50 2812.50 + 500 = 3312.50 3312.50 = 3312.50 (correct) adult tickets (x) sole were 375 child tickets (y) sold were 125

OpenStudy (anonymous):

THANKS SO MUCH!

OpenStudy (texaschic101):

no problem...sorry it took so long...I messed up and had to start again, but I finally got it :)

OpenStudy (anonymous):

:D